# From Distance to Angle: One-Shot Detection Under Additive White Cauchy Noise

### 10 Apr 2026 v1

@@ V Conclusion
This paper developed a geometric asymptotic perspective on one-shot detection under additive white Cauchy noise. Although isotropic Cauchy noise induces the same Euclidean Voronoi decision regions as Gaussian noise, the corresponding reliability laws differ fundamentally because the probability mass assigned to those regions is redistributed by the heavy-tailed noise geometry. In the small-noise regime, we established a reciprocal distance-spectrum upper bound showing that error probabilities retain sensitivity to the full constellation geometry. In the large-noise regime, we proved that the correct-decision probability converges to an angular limit determined solely by the Voronoi recession cone. Together, these results formalize a regime-dependent transition from distance-based to angle-based reliability descriptors. The worked examples in Appendices A and B show that this transition can lead either to highly nonuniform symbol reliability, including complete geometric collapse, or to symmetric limiting behavior in standard constellations such as 4QAM.
@@ Appendix A Geometric Collapse: A Four-Point Example
yet their large-noise behavior under AWCN differs substantially because their Voronoi recession cones have different angular measures.
[…]
Figure 1: Voronoi geometry for the constellation $\mathcal{X}$ (see Appendix A). Left: Euclidean Voronoi cells. Right: recession cones and their angular patches on $\mathbb{S}^{1}$. The point $P_{1}$ has a degenerate cone of zero angular measure, whereas $P_{2},P_{3},P_{4}$ retain nontrivial patches. The nondegenerate patches correspond to $P_{2}$, $P_{3}$, and $P_{4}$, with opening angles $\pi/2$, $3\pi/4$, and $3\pi/4$, respectively.
The AWGN column is included only as a qualitative comparison: under Gaussian noise, the nearest-neighbor structure largely governs the high-reliability regime, whereas under AWCN the small-noise upper bound depends on the full reciprocal distance spectrum and the large-noise limit depends only on angular measures.
@@ A-A Small-Noise Limit: Explicit Distance-Spectrum Evaluation
These calculations reproduce exactly the AWCN small-$\gamma$ column in Table I. In particular, although all four points share the same minimum distance, their reciprocal distance sums differ, so Theorem 1 already distinguishes them at leading order.
TABLE I: Quantitative transition from distance-based to angular descriptors for $\mathcal{X}$ (an illustrative example in Appendix A).
Constellation Point AWGN AWCN AWCN
$x$ (small $\sigma$) (small $\gamma$) ($\gamma\to\infty$)

### 20 Apr 2026 v2

@@ VI Conclusion
This paper developed a geometric asymptotic perspective on one-shot detection under isotropic multivariate Cauchy noise. Although isotropic Cauchy noise induces the same Euclidean Voronoi decision regions as Gaussian noise, the corresponding reliability laws differ fundamentally because the probability mass assigned to those regions is redistributed by the heavy-tailed noise geometry. In the small-noise regime, we established a reciprocal distance-spectrum upper bound showing that error probabilities retain sensitivity to the full constellation geometry. In the large-noise regime, we proved that the correct-decision probability converges to an angular limit determined solely by the Voronoi recession cone. Together, these results formalize a regime-dependent transition from distance-based to angle-based reliability descriptors. The worked examples in Appendices A and B, together with the finite-$\gamma$ numerical validations in Section V, show that this transition can lead either to highly nonuniform symbol reliability, including complete geometric collapse, or to symmetric limiting behavior in standard constellations such as 4QAM.
@@ Appendix A Geometric Collapse: A Four-Point Example
yet their large-noise behavior under isotropic Cauchy noise differs substantially because their Voronoi recession cones have different angular measures.
The AWGN column is included only as a qualitative comparison: under Gaussian noise, the nearest-neighbor structure largely governs the high-reliability regime, whereas under the isotropic Cauchy model the small-noise upper bound depends on the full reciprocal distance spectrum and the large-noise limit depends only on angular measures.
@@ A-A Small-Noise Limit: Explicit Distance-Spectrum Evaluation
These calculations reproduce exactly the isotropic Cauchy small-$\gamma$ column in Fig. 3(b). In particular, although all four points share the same minimum distance, their reciprocal distance sums differ, so Theorem 1 already distinguishes them at leading order.
[…]
(a) Geometric properties
Constellation Point AWGN Isotropic Cauchy Isotropic Cauchy
$x$ (small $\sigma$) (small $\gamma$) ($\gamma\to\infty$, $P_{e}$)
@@
(b) Quantitative reliability descriptors
Figure 3: Geometric and quantitative analysis for the illustrative four-point constellation $\mathcal{X}$ (see Appendix A). (a) Left: Euclidean Voronoi cells. Right: recession cones and their angular patches on $\mathbb{S}^{1}$. The point $P_{1}$ has a degenerate cone of zero angular measure, whereas $P_{2},P_{3},P_{4}$ retain nontrivial patches with opening angles $\pi/2$, $3\pi/4$, and $3\pi/4$, respectively. (b) The corresponding theoretical reliability descriptors showcasing the quantitative transition from Gaussian minimum-distance behavior to Cauchy distance-spectrum and angular limits.